OpenStudy (anonymous):

8y - 1 = x 3x = 2y Solve the system of equations by substitution. a.(1/33, 1/11) b.(1/11, 3/22) c.no solution d.coincident

OpenStudy (anonymous):

3 [ 8y -1] = 2y 24y - 3 = 2y 22y = 3 => y = 3/22 substitute value of y in , 8y-1=x 8(3/22) -1 = x => x = 12/11 - 1 x = 1/11 So, (b)

OpenStudy (anonymous):

what about 7 + 2y = 8x 3x - 2y = 0 Solve the system of equations by substitution. (70/3, 140/9) (7/5, 21/10) no solution coincident